OpenStudy (anonymous):

Solve: −2(x − 8) = −10

OpenStudy (anonymous):

Hint: Distribute the -2. Then solve for x

OpenStudy (anonymous):

divide both sides by -2

OpenStudy (anonymous):

3

OpenStudy (anonymous):

That's easier^

OpenStudy (anonymous):

or -2x +16 =-10 -2x = -10-16 -2x = 26 -x=13 x=13